# Reward Is Enough: LLMs Are In-Context Reinforcement Learners

### v1

@@ Abstract
Reinforcement learning (RL) is a human-designed framework for solving sequential decision making problems. In this work, we demonstrate that, surprisingly, RL emerges in LLM’s (Large Language Model) inference time – a phenomenon known as in-context RL (ICRL). Specifically, we propose a novel multi-round prompting framework called ICRL prompting. The goal is to prompt the LLM to complete a task. After the LLM generates a response at the current round, we give numerical scalar feedbacks for the response, called the rewards. At the next round, we prompt the LLM again with the same task and a context consisting of all previous responses and rewards. We observe that the quality of the LLM’s response increases as the context grows. In other words, the LLM is able to maximize the scalar reward signal in the inference time, just like an RL algorithm. We evaluate ICRL prompting in three benchmarks (Game of 24, creative writing, and ScienceWorld) and demonstrate significant performance improvements over baseline methods such as Self-Refine and Reflexion. Surprisingly, in some experiments the reward signals are generated by the LLM itself, yet performance improvements are still observed from ICRL prompting, offering a promising paradigm for scaling test-time compute.
@@ 1 Introduction
For Large Language Models (LLMs) to act as effective agents to solve novel tasks, the capacity for LLMs to improve during the inference time is essential. Learning and search are the two general methods that can leverage scaling computation for performance improvement (Sutton,, 2019). Search has been successfully applied to LLM’s self-improvement during the inference time, starting from the simple Best-of-N (Stiennon et al.,, 2022) to Tree of Thoughts (Yao et al., 2023a, ) and Monte Carlo Tree Search (Ding et al.,, 2024). Learning, however, has yet to receive the same attention for LLM’s self-improvement during the inference time. In-context (supervised) learning (ICL, Brown et al., (2020)), as a supervised learning paradigm during the inference time, cannot bring performance higher than the (expert) demonstrations in the context. Therefore, ICL does not allow the LLM to learn from its own experience (Silver and Sutton,, 2024) to self-improve.
Reinforcement Learning (RL, Sutton and Barto, (2018)) is perhaps the most successful human designed framework for sequential decision making. In this work, we demonstrate that, surprisingly, RL also emerges in LLM’s inference time. Such inference-time reinforcement learning process is commonly called in-context reinforcement learning (ICRL, Laskin et al., (2023); Moeini et al., (2025)). In this paper, we propose a multi-round prompting framework called ICRL prompting, enabling LLM’s self-improvement during the inference time through (reinforcement) learning. The goal of ICRL prompting is to prompt the LLM to complete a task. Initially, the prompt is only the task description. After the LLM generates a response, we give numerical scalar feedbacks for the response, called the rewards. Then in the next round, we prompt the LLM again with the same task description and a context consisting of all previous responses and rewards. So on and so forth. We observe that the quality of the LLM’s response increases as the context grows. In other words, the LLM is able to maximize the scalar reward signal during the inference time, just like an RL algorithm. This demonstrates the potential of learning methods for LLM’s self-improvement during the inference time. Notably, the scalar reward signal is the only feedback we provide to the LLM. This complies with both the reward hypothesis (Sutton,, 2004), “that all of what we mean by goals and purposes can be well thought of as maximization of the expected value of the cumulative sum of a received scalar signal (reward)”, and the “reward is enough” hypothesis (Silver et al.,, 2021), “intelligence, and its associated abilities, can be understood as subserving the maximisation of reward”.
@@
(1): We propose the ICRL prompting framework that demonstrates LLM’s inference-time self-improvement through RL. One key technical challenge is designing this framework to be as minimal as possible, ensuring that the observed improvements clearly stem from the LLM’s intrinsic capacity of ICRL, rather than from auxiliary external code or additional engineered LLM interactions. Specifically, our framework avoids the use of techniques such as alternating different sampling temperatures (Zhang et al.,, 2024), extra sampling steps (Yang et al.,, 2024), any form of prioritized experience replay (e.g., subsampling, selecting or re-ordering past trials based on rewards), textual descriptions of reward or textual gradients (Yuksekgonul et al.,, 2025) from additional LLM nodes. By adhering to this minimal design, we provide strong evidence that by merely providing instructions on exploration and exploitation, we can elicit the ICRL capabilities from the pretrained LLM itself.
(2): We provide strong evidence suggesting the emergence of RL in LLM’s inference time when the ICRL prompting framework is used. Specifically, we demonstrate the maximisation of the scalar reward signal, the exploration-exploitation trade-off in LLM’s inference time, the performance improvements from the growth of the context, the performance drop with short context, and the performance drop when the reward is absent. All those observations are well expected for an RL algorithm. Essentially, this is a “duck test” (Heim,, 2007) [^1] If it looks like a duck, swims like a duck, and quacks like a duck, then it probably is a duck. for the inference process.
(3): We demonstrate significant performance improvements of ICLR prompting over baseline methods like Self-Refine (Madaan et al.,, 2023) and Reflexion (Shinn et al.,, 2023), in three different benchmarks including Game of 24, creative writing, and ScienceWorld. In Game of 24 and creative writing, the scalar reward signal is generated by the LLM itself. Yet performance improvements are still observed.
@@ 2 Background
Reinforcement Learning. RL uses Markov Decision Processes (MDPs) to model a task, consisting of a state space $\mathcal{S}$, an action space $\mathcal{A}$, a reward function $r:\mathcal{S}\to\mathbb{R}$, an initial distribution $p_{0}\in\Delta(\mathcal{S})$ with $\Delta(\mathcal{S})$ denoting the set of probability distributions over $\mathcal{S}$, and a transition function $p:\mathcal{S}\times\mathcal{A}\to\Delta(\mathcal{S})$. At time step 0, an initial state $S_{0}$ sampled from $p_{0}$. At time $t$, an agent at $S_{t}$ takes an action $A_{t}$ according to its policy $\pi:\mathcal{S}\to\Delta(\mathcal{A})$ with $\Delta(\mathcal{A})$ denoting the set of probability distributions over $\mathcal{A}$, i.e., $A_{t}\sim\pi(S_{t})$. The action $A_{t}$ is then executed, after which the agent transitions to a successor state $S_{t+1}\sim p(S_{t},A_{t})$ and recieves a reward $R_{t+1}\doteq r(S_{t+1})$. This agent-environment interaction continues until a time $T$, which marks the end of an episode. The goal of the agent is to adapt its policy $\pi$ such that the expected total rewards $J(\pi)\doteq\mathbb{E}[\sum_{t=1}^{T}R_{t}]$ is maximized. In modern deep RL (Mnih et al.,, 2015; Schulman et al.,, 2017), the policy $\pi$ is usually parameterized by a neural network. We use $\theta$ to denote the network parameter and write the policy as $\pi_{\theta}$. Typically, RL algorithms update $\theta$ to adapt its policy. For example, at time $t$, the action $A_{t}$ is sampled from $\pi_{\theta_{t}}(S_{t})$. The RL algorithm then update $\theta_{t}$ to $\theta_{t+1}$ based on available information such as $S_{0},A_{0},R_{1},\dots,S_{t},A_{t},R_{t+1},S_{t+1}$. Then at time $t+1$, the action $A_{t+1}$ is sampled from the updated policy $\pi_{\theta_{t+1}}(S_{t+1})$. Essentially, the RL process is reflected in the udpates of $\quantity{\theta_{t}}$ and we call this in-weight RL.
@@
In-Context Reinforcement Learning. ICRL (Moeini et al.,, 2025), first coined by Laskin et al., (2023), is an emerging inference-time compute paradigm where the RL process occurs in the inference time (i.e., the forward pass) of the network without any parameter update. In ICRL, the policy $\pi_{\theta}$ is additionally conditioned on a context called $C_{t}$, i.e., $A_{t}\sim\pi_{\theta}(S_{t},C_{t})$. The construction of $C_{t}$ is an active research area but one example is to use all previous state-action-reward pairs obtained in the task. Notably, this usually includes state-action-reward pairs from all previous episodes, not just the current episode (Laskin et al.,, 2023). In ICRL, there is a pretraining stage where the network $\theta$ is pretrained in a wide range of tasks (MDPs). We use $\theta_{*}$ to denote the parameter after the pretraining. After the pretraining stage, the policy $\pi_{\theta_{*}}$ is evaluated in new tasks. In other words, in the new MDP, the action $A_{t}$ is sampled from $\pi_{\theta_{*}}(S_{t},C_{t})$. Importantly, the $\theta_{*}$ is kept fixed. Nevertheless, it is observed that the quality of $A_{t}$ increases as $C_{t}$ grows in the new task. Since $\theta_{*}$ is fixed, this improvement can only from the increase of the context. This is thus called in-context policy improvement. Notably, this in-context policy improvement is also observed even when the new task is out of the distribution of the pretraining tasks, e.g., Laskin et al., (2023) demonstrate in-context policy improvement in new bandit problems that have the opposite optimal arms to the pretraining bandit problems. Thus this in-context policy improvement cannot be attributed to the hypothesis that $\theta_{*}$ memorizes the pretraining tasks. The only plausible hypothesis seems to be that the forward pass of the network parameterized by $\theta_{*}$ implements some RL algorithm to process the information in the context $C_{t}$ to generate the action $A_{t}$. This inference-time (forward pass) RL is called in-context RL.
@@
LLMs. The token generation process of LLMs can be modeled via RL. In short, the state is all generated tokens and the action is the next token to generate. Namely, let $\mathcal{V}$ be the set of all possible tokens. We consider a state space $\mathcal{S}\doteq\bigcup_{i=1}^{\infty}\mathcal{V}^{i}$ and an action space $\mathcal{A}\doteq\mathcal{V}$. At time step 0, an initial prompt is given, denoted as $S_{0}\in\mathcal{S}$. In this work, $S_{0}$ contains a description of a task. We refer to the LLM with parameter $\theta$ as $\pi_{\theta}$. At time $t$, given the current tokens $S_{t}$, a new token $A_{t}$ is sampled from $\pi_{\theta}(S_{t})$. The new state is then $S_{t+1}=\matrixquantity[S_{t}A_{t}]$, i.e., the new state is obtained by concatenating current tokens and the new token. A reward signal $R_{t+1}\doteq r(S_{t+1})$ is then emitted according to a reward funciton $r$. This token generation process continues until a time $T$, where either $T$ is the maximal allowed response length or $A_{T-1}$ is a special end-of-sequence token. Either way, this marks the end of an episode and the final state $S_{T}$, called the terminal state, contains both the initial task description and the LLM’s response. There are two types of reward functions. One is sparse (the outcome reward model, Ouyang et al., (2022)), where $r(s)$ is nonzero only when $s$ is a terminal state. The other is dense (the progress reward model, Lightman et al., (2023)), where $r(s)$ can also be nonzero for non-terminal states.
@@ 3 In-Context Reinforcement Learning Prompting
7:     // $[A_{0}\,A_{1},\dots,A_{T-1}]$ is the LLM’s response to $s_{\text{task}}$ at the current episode
@@ 4 Related Works
The study of inference time RL algorithms dates back to Duan et al., (2016); Wang et al., (2016) and recently Laskin et al., (2023) coined the word “in-context reinforcement learning”, after which the interest in ICRL grew quickly (Kirsch et al.,, 2023; Raparthy et al.,, 2023; Schmied et al.,, 2024; Lee et al.,, 2024; Zisman et al.,, 2023; Grigsby et al.,, 2024; Lu et al.,, 2023; Bauer et al.,, 2023; Wang et al.,, 2025; Cook et al.,, 2024; Xu et al.,, 2024; Shi et al.,, 2024; Huang et al.,, 2024; Liu and Abbeel,, 2023; Dai et al.,, 2024) See Moeini et al., (2025) for a comprehensive survey. Most existing ICRL works are essentially a subarea of meta RL (Beck et al.,, 2023). They use games or robotics as benchmark and the network size is small (compared with pretrained LLMs like GPT-4.1). They focus on algorithms to induce this ICRL capability and train sequence models from scratch. By contrast, this work studies the ICRL capability of standard pretrained LLMs without any parameter updates. Nevertheless, some existing ICRL works do include pretrained LLM. For example, Brooks et al., (2024) use a pretrained LLM as a world model to simulate rollouts; Mirchandani et al., (2023) treated LLMs as general sequence predictors. Closest works to ours study LLMs in bandit tasks, specifically Multi-Arm Bandits (MAB) (Krishnamurthy et al.,, 2024; Nie et al.,, 2024; Park et al.,, 2024) and Contextual Bandits (CB) (Nie et al.,, 2024; Monea et al.,, 2025). They show that artificial interventions, such as including statistics commonly used in MAB algorithms in the context, can improve LLM’s performance on various MAB and CB tasks. In particular, Monea et al., (2025) reformulate a classification task as a CB problem by assigning a binary reward based on classification correctness. Their comprehensive empirical results indicate that LLMs perform poorly on this task without either prompts that encourage exploration or additional mechanisms to filter incorrect actions from the context. Overall, these works demonstrate that LLMs remain uncompetitive with algorithmic baselines for both CB and MAB, even after such interventions. We shift our focus away from tasks solvable by classic analytic bandit algorithms and instead tackle text tasks that require real-world knowledge about math, science, and linguistics, domains where LLMs uniquely excel. Our experiments show that interaction history alone, when paired with numerical rewards, outperforms baseline test-time self-improvement methods, including various prompting and sampling techniques. Importantly, we do not filter any “failed” action with low rewards since a key characterization of RL is to learn from failure as well.
Broadly speaking, ICRL can be viewed as a subarea of in-context learning (ICL, Brown et al., (2020)) if ICL is taken to encompass any inference-time learning paradigm. In practice, however, ICL often refers specifically to in-context supervised learning (ICSL) (Brown et al.,, 2020). The key difference between ICRL and ICSL is that in ICSL, the context consists of task-demonstration pairs, where demonstrations are typically generated by an expert. In contrast, ICRL contexts contain tasks, previous attempts, and associated rewards. Agarwal et al., (2024) show that scaling ICL with synthetic answers can outperform few-shot expert ICL, but their rule-based verification assumes access to many auxiliary problems drawn from the same distribution as the current task. In our setting we interact only with the current task and receive either a synthetic or a ground truth scalar reward, without access to expert or successful trajectories. Like any reinforcement learning method, ICRL exploits both rewards and trajectories, allowing the model to improve on the current task by learning from all of its prior failures.

### v3

@@ Abstract
Reinforcement learning (RL) is a framework for solving sequential decision-making problems. In this work, we demonstrate that, surprisingly, RL emerges during the inference time of large language models (LLMs), a phenomenon we term in-context RL (ICRL). To reveal this capability, we introduce a simple multi-round prompting framework, we call ICRL prompting, for inference-time self-improvement. The goal of ICRL prompting is to guide LLMs to perform reinforcement learning during inference for self-improvement on a given task. After each response, the model receives numerical scalar feedback, denoted as a reward. In the next round, we prompt the LLM again together with a context that concatenates all prior responses and their associated rewards. We consistently observe that response quality improves as the context grows. In other words, the LLM can optimize scalar reward signals during inference, exhibiting behavior analogous to reinforcement learning. We evaluate ICRL prompting on Game of 24, creative writing, ScienceWorld, and Olympiad-level math competitions (AIME and HMMT), demonstrating significant improvements over baselines such as Self-Refine and Reflexion. Notably, even when the reward signals are generated by the same LLM, ICRL prompting still improves performance, highlighting a promising new paradigm for test-time scaling.
22footnotetext: Equal contribution.
@@ 1 Introduction
For Large Language Models (LLMs) to act as effective agents on novel tasks, they must be able to improve during inference time, a capability often referred to as test-time scaling (tts). Learning and search are the two general methods that can leverage scaling computation for performance improvement (sutton2019bitter), reaching superhuman performance on Chess (campbell2002deepblue) and Go (silver2016mastering). Search has been successfully applied to LLM self-improvement in test-time scaling, starting from the simple Best-of-N (best-of-n) to Tree of Thoughts (tot) and Monte Carlo Tree Search (everything-of-thought).
Learning, however, has yet to receive the same attention for LLM self-improvement at inference time. In-context (supervised) learning (ICL; ICL), as a supervised learning paradigm, requires expert demonstrations as ground-truth labels. However, such demonstration data are not easily scalable during inference time, which restricts the applicability of ICL to test-time scaling. Thus, LLMs must instead learn from their own generated experience for continual self-improvement (silver2024era).
Reinforcement learning is perhaps the most successful algorithm capable of self-improvement independent of human knowledge (silver2017mastering). However, its major successes have primarily appeared in simulated environments (mnih2015human; silver2016mastering) or during the training time of LLMs guo2025deepseek. These current RL settings fall short in the big world setting (javed2024bigworld), where the real-world environment is vastly more complex than the agent itself. In such environments, agents will encounter numerous situations far beyond their prior training data and must adapt and improve their solutions on the fly. Bridging this gap requires models that (1) can handle diverse tasks in the real world, where natural language often constitutes an essential action space (silver2025welcome), and (2) can continually improve their solutions during inference, rather than relying on costly retraining for every novel situation.
This naturally raises the question: can reinforcement learning emerge during the inference phase of LLMs? Enabling LLMs to perform RL purely in context provides an elegant mechanism to meet both requirements: LLM provides a general-purpose initial policy, while RL introduces the capability for continual self-improvement. Inspired by the first surprising evidence that LLMs can act as in-context learners in supervised settings (ICL), a growing body of work has begun to explore in-context reinforcement learning (ICRL; moeini2025survey). However, current instantiations are largely restricted to bandit or simulated environments (monea2025llmsincontextbanditreinforcement; krishnamurthy2024can), failing short of addressing many diverse open-ended tasks where natural language is the action space.
In this paper, we bridge this critical gap by demonstrating that LLMs can act as effective in-context reinforcement learners, an emergent capability that improves performance on diverse, language-based tasks ranging from conducting scientific experiments to creative writing to solving olympiad-level mathematics. To reveal this capability, we introduce a simple multi-round prompting framework, ICRL prompting. The goal of ICRL prompting is to guide LLMs to perform reinforcement learning for self-improvement on a task. Initially, the prompt is only the task description. After the LLM generates a response, we give numerical scalar feedbacks for the response, called the rewards. Then in the next round, we prompt the LLM again with the same task description and a context consisting of all previous responses and rewards. So on and so forth. We observe that the quality of the LLM’s response increases as the context grows. In other words, the LLM is able to maximize the scalar reward signal during the inference time, just like an RL algorithm.
A key design principle of ICRL prompting is minimality. To ensure that the observed gains arise from the emergent RL capacity of LLMs rather than auxiliary mechanisms, we deliberately exclude textual gradients (textgrad), prioritized experience replay, sampling-based heuristics (zhang2024edtimprovinglargelanguage; llm_as_optimizers), or additional engineered modules (brooks2024large). The only supervision provided is the scalar reward itself. This complies with both the reward hypothesis (reward-hypothesis), “that all of what we mean by goals and purposes can be well thought of as maximization of the expected value of the cumulative sum of a received scalar signal (reward)”, and the “reward is enough” hypothesis (reward-is-enough), “intelligence, and its associated abilities, can be understood as subserving the maximisation of reward”.
@@
(1): We introduce the ICRL prompting framework, a minimal design that elicits inference-time self-improvement in LLMs using only scalar rewards. Just as ICL places $(x,y)$ pairs in context, our framework places state–action–reward tuples with simple meta-instructions in context. This design isolates the LLM’s intrinsic capacity for ICRL, free from external code or engineered mechanisms.
(2): We provide strong evidence suggesting the emergence of RL in LLM’s inference time when the ICRL prompting framework is used. Specifically, we demonstrate the maximisation of the scalar reward signal, the exploration-exploitation trade-off in LLM’s inference time, the performance improvements from the growth of the context, the performance drop with short context, and the performance drop when the reward is absent. All those observations are well expected for an RL algorithm. Essentially, this is a “duck test” (heim2007exploringindiana) [^1] If it looks like a duck, swims like a duck, and quacks like a duck, then it probably is a duck. for the inference process.
(3): We demonstrate that ICRL prompting yields significant improvements over self-revision methods such as Self-Refine (self-refine) and Reflexion (shinn2023reflexion), across diverse benchmarks including Game of 24, creative writing, ScienceWorld, and Olympiad-level mathematics (AIME and HMMT). In Game of 24 and creative writing, the rewards are generated by the LLM itself, yet consistent performance gains are still observed.
@@ 2 Background
Reinforcement Learning. RL uses Markov Decision Processes (MDPs) to model a task, consisting of a state space $\mathcal{S}$, an action space $\mathcal{A}$, a reward function $r:\mathcal{S}\to\mathbb{R}$, an initial distribution $p_{0}\in\Delta(\mathcal{S})$ with $\Delta(\mathcal{S})$ denoting the set of probability distributions over $\mathcal{S}$, and a transition function $p:\mathcal{S}\times\mathcal{A}\to\Delta(\mathcal{S})$. At time step 0, an initial state $S_{0}$ sampled from $p_{0}$. At time $t$, an agent at $S_{t}$ takes an action $A_{t}$ according to its policy $\pi:\mathcal{S}\to\Delta(\mathcal{A})$ with $\Delta(\mathcal{A})$ denoting the set of probability distributions over $\mathcal{A}$, i.e., $A_{t}\sim\pi(S_{t})$. The action $A_{t}$ is then executed, after which the agent transitions to a successor state $S_{t+1}\sim p(S_{t},A_{t})$ and recieves a reward $R_{t+1}\doteq r(S_{t+1})$. This agent-environment interaction continues until a time $T$, which marks the end of an episode. The goal of the agent is to adapt its policy $\pi$ such that the expected total rewards $J(\pi)\doteq\mathbb{E}[\sum_{t=1}^{T}R_{t}]$ is maximized. In modern deep RL (mnih2015human; schulman2017proximal), the policy $\pi$ is usually parameterized by a neural network. We use $\theta$ to denote the network parameter and write the policy as $\pi_{\theta}$. Typically, RL algorithms update $\theta$ to adapt its policy. For example, at time $t$, the action $A_{t}$ is sampled from $\pi_{\theta_{t}}(S_{t})$. The RL algorithm then update $\theta_{t}$ to $\theta_{t+1}$ based on available information such as $S_{0},A_{0},R_{1},\dots,S_{t},A_{t},R_{t+1},S_{t+1}$. Then at time $t+1$, the action $A_{t+1}$ is sampled from the updated policy $\pi_{\theta_{t+1}}(S_{t+1})$. Essentially, the typical RL process is reflected in the updates of $\theta_{t}$}.
@@
In-Context Reinforcement Learning. ICRL (moeini2025survey), first coined by laskin2023incontext, is an emerging inference-time compute paradigm where the RL process occurs in the inference time (i.e., the forward pass) of the network without any parameter update. In ICRL, the policy $\pi_{\theta}$ is additionally conditioned on a context called $C_{t}$, i.e., $A_{t}\sim\pi_{\theta}(S_{t},C_{t})$. The construction of $C_{t}$ is an active research area but one example is to use all previous state-action-reward pairs obtained in the task. Notably, this usually includes state-action-reward pairs from all previous episodes, not just the current episode (laskin2023incontext). In ICRL, there is a pretraining stage where the network $\theta$ is pretrained in a wide range of tasks (MDPs). We use $\theta_{*}$ to denote the parameter after the pretraining. After the pretraining stage, the policy $\pi_{\theta_{*}}$ is evaluated in new tasks. In other words, in the new MDP, the action $A_{t}$ is sampled from $\pi_{\theta_{*}}(S_{t},C_{t})$. Importantly, the $\theta_{*}$ is kept fixed. Nevertheless, it is observed that the quality of $A_{t}$ increases as $C_{t}$ grows in the new task. Since $\theta_{*}$ is fixed, this improvement can only from the increase of the context. This is thus called in-context policy improvement. Notably, this in-context policy improvement is also observed even when the new task is out of the distribution of the pretraining tasks, e.g., laskin2023incontext demonstrate in-context policy improvement in new bandit problems that have the opposite optimal arms to the pretraining bandit problems. Thus this in-context policy improvement cannot be attributed to the hypothesis that $\theta_{*}$ memorizes the pretraining tasks. The only plausible hypothesis seems to be that the forward pass of the network parameterized by $\theta_{*}$ implements some RL algorithm to process the information in the context $C_{t}$ to generate the action $A_{t}$. This inference-time (forward pass) RL is called in-context RL.
@@
LLMs as RL Agents. The token generation process of LLMs can be modeled via RL. In short, the state is all generated tokens and the action is the next token to generate. Namely, let $\mathcal{V}$ be the set of all possible tokens. We consider a state space $\mathcal{S}\doteq\bigcup_{i=1}^{\infty}\mathcal{V}^{i}$ and an action space $\mathcal{A}\doteq\mathcal{V}$. At time step 0, an initial prompt is given, denoted as $S_{0}\in\mathcal{S}$. In this work, $S_{0}$ contains a description of a task. We refer to the LLM with parameter $\theta$ as $\pi_{\theta}$. At time $t$, given the current tokens $S_{t}$, a new token $A_{t}$ is sampled from $\pi_{\theta}(S_{t})$. The new state is then $S_{t+1}=\matrixquantity[S_{t}A_{t}]$, i.e., the new state is obtained by concatenating current tokens and the new token. A reward signal $R_{t+1}\doteq r(S_{t+1})$ is then emitted according to a reward function $r$. This token generation process continues until a time $T$, where either $T$ is the maximal allowed response length or $A_{T-1}$ is a special end-of-sequence token. Either way, this marks the end of an episode and the final state $S_{T}$, called the terminal state, contains both the initial task description and the LLM’s response. There are two types of reward functions. One is sparse (the outcome reward model, ouyang2022training), where $r(s)$ is nonzero only when $s$ is a terminal state. The other is dense (the progress reward model, lightman2023lets), where $r(s)$ can also be nonzero for non-terminal states.
@@ 3 In-Context Reinforcement Learning Prompting
7:  // $A_{0}\,A_{1},\dots,A_{T-1}]$ is the LLM’s response to $s_{\text{task}}$ at the current episode
@@ 4 Related Works
The study of inference-time RL algorithms dates back to duan2016rl; wang2016learning, with laskin2023incontext later coining the term in-context reinforcement learning (ICRL), spurring rapid growth in the field (kirsch2023towards; raparthy2023generalization; schmiedLargeRecurrentAction2024; lee2024supervised; zisman2023emergence; grigsby2023amago; lus42023; bauer2023human; wang2025transformers; cook2024artificial; xu2024meta; shi2024cross; huang2024context; liuEmergentAgenticTransformer2023; dai2024context). See moeini2025survey for a survey. Most existing ICRL works, as a subarea of meta-RL (beck2023survey), use small models trained from scratch in games or robotics. Some employ pretrained LLMs, e.g., as simulators (brooks2024large; mirchandani_large_2023; resendiz2025parlpromptbasedagentsreinforcement) or in bandit tasks (krishnamurthy2024can; nieEVOLvEEvaluatingOptimizing2024; park2024llm; monea2025llmsincontextbanditreinforcement), where artificial interventions are often needed and LLMs remain uncompetitive with algorithmic baselines.